OpenStudy (anonymous):

Stuck! Help Me! Solve the following system using the substitution method. 16x + 3y = 0 8x + y = 1

OpenStudy (anonymous):

\[(1)\quad 16x + 3y = 0 \\ (2)\quad 8x + y = 1\] From 2 you can see that \(\Large y = 1 - 8x\), substitute that in 1 to get: \[16x + 3(1 - 8x) = 0\\ 16x + 3 - 24x = 0\\ -8x = -3\\ x = \frac{3}{8}\] then substitute x = 3/8 in any of the equations to find y: \[8\left(\frac{3}{8}\right)+y = 1\\ y = -2\] Then the solution is \(\Large (x, y ) = \left(\frac{3}{8}, -2\right)\)

OpenStudy (anonymous):

Thank you :)